# CLEAR: Character Unlearning in Textual and Visual Modalities

### v1

@@ 4.3 Evaluation Metrics
5 Experiments
First, we explore the capabilities of the current unlearning methods within single domains in Sec. 5.1 and 5.2. Second, we transfer them to textual-visual MMU in Sec. 5.3.
5.1 Unlearning Textual Domain (LLMs)
For the experiments on the textual domain exclusively, we consider the textual part of the proposed CLEAR dataset and use the LLMs, which are often used in MLLMs, specifically Llama2-7B Touvron et al. (2023) and Mistral-7B Jiang et al. (2023). First, we finetune the model on all the data and unlearn it on the forget set. To evaluate the final quality of unlearning, we calculate Real, Retain, and Forget metrics and Forget Quality. The full results are provided in Table 1. See the Appendix D for pipeline details and hyperparameters.
[…]
Original —– 0.48 0.3 0.51 -61.22
Gold —– 0.50 0.19 0.51 0.00
LLMU text 0.47 0.37 0.49 -71.23
LLMU visual 0.50 0.35 0.51 -60.26
LLMU both 0.47 0.25 0.51 -95.12
SCRUB text 0.49 0.35 0.51 -61.22
SCRUB visual 0.48 0.37 0.49 -60.26
SCRUB both 0.49 0.36 0.52 -60.26
DPO text 0.46 0.38 0.49 -62.18
DPO visual 0.49 0.22 0.49 -90.26
DPO both 0.46 0.22 0.48 -91.46
Table 3: Results of unlearning of different modalities. We finetune on full datasets (both modalities), then forget on a single domain subset (text or visual) or full forget set. Original – model before unlearning. Gold - a model trained only on retain.
5.2 Unlearning Visual Domain
Experiments on the visual domain of the CLEAR dataset focus on the face biometrics (identification) task. While face identification can be treated as a classification task with a fixed set of individuals, it is typically framed as a few-shot or metric learning problem, aiming to train an embedding model that maps images of the same person to nearby points in the embedding space and ensures separation between different individuals and this should hold even for the unseen during training persons. We chose ResNet-18 He et al. (2015) due to its relatively small size and scalability.
We fine-tuned ResNet-18, pre-trained on ImageNet, for 100 epochs on the Celebs dataset Zhang et al. (2020), using photos of 797 individuals for training and 200 for testing. To compute identification accuracy, we averaged the embeddings of five images per individual to obtain reference (enrollment) vectors and classified the remaining images using cosine similarity. The model achieved 99.0% accuracy on the training set and 83.4% on the test set. For unlearning evaluation, the model was fine-tuned on the visual part of CLEAR dataset 128 times, following a Membership Inference Attack (MIA) approach. We trained 64 models with the forget included and 64 models without it, alternating with the holdout set (if the forgetting set is used for training, the holdout set is not, and vice versa).
We evaluated the models on $D_{F}$, $D_{R}$, and $D_{H}$ using accuracy, U-LIRA Hayes et al. (2024), and U-MIA attack metrics. U-MIA is a lightweight, population-based attack that trains a binary classifier on shadow model outputs and their status (1 if from the forget set, 0 otherwise). Successful unlearning occurs when Holdout Accuracy closely aligns with Forget Accuracy. From the standpoint of the attack, U-LIRA and U-MIA should be unable to distinguish between the two sets, resulting in an attack accuracy close to 0.50. This situation would indicate that the model has "forgotten" the specific data, as the model treats the forgetting and holdout sets similarly. We compare the forget and holdout sets logit distributions in Appendix 6.
5.3 Multimodal Experiments
For the source model, we use LLaVa model Liu et al. (2023) with ViT Dosovitskiy et al. (2021) as visual encoder and LLaMa2-7B Touvron et al. (2023) as a language model. First, we finetune it on the full CLEAR, both visual and textual parts, and call this model "original", as it contains forget and retain sets of knowledge. Then, we perform the unlearning process on it. We use the same hyperparameters for each method. Then, we evaluate the unlearned model according to our metric setup described in Sec. 4.3. For comparison, we demonstrate the metrics of the "gold" model. The results of experiments and corresponding metrics are provided in Table 4. Details of the experiments pipeline are described in Appendix F.
5.3.1 Is Textual Unlearning Enough?
We begin by addressing the question: Can we forget a person using only textual data, and does multimodality introduce new challenges to unlearning?
To explore this, we attempt to forget 20 individuals from the forget set using only textual unlearning. We also perform unlearning using only visual data or both modalities, and then compare the results. We find that unlearning text alone is sufficient to achieve a low forget metric, consistent with previous findings. However, this also results in a noticeable drop in retain metrics. Full results are provided in Table 3.
5.3.2 Unlearning Both Domains
After we understand that multimodal unlearning can not be fully addressed using a single modality, we proceed with experiments on unlearning across both modalities. We take our source model $f_{\theta}$ and apply unlearning methods. As forget set, we use all the available data about 20 persons - 10% of the dataset size.
5.4 LoRA Regularization
As shown previously Jia et al. (2024), model sparsity can improve unlearning, but significant drops in retain metrics still occur. We hypothesize that keeping the model close to its initial state during unlearning could help preserve retain knowledge.
LoRA adapters Hu et al. (2021) have become a standard technique to reduce computational demands in large-scale NLP models. We propose using the magnitude of LoRA weights as a proxy for how far the model has deviated from its initial state. To address this, we add the $\ell_{1}$ norm of the adapter weights to the unlearning loss, using a fixed $\lambda=0.01$ without tuning, though tuning could improve results. The results of the experiments are shown in Table 4.
Method LoRA L1 Regularization Real $\uparrow$ Retain metric $\uparrow$ Forget metric $\downarrow$ Log Forget Quality $\uparrow$
Original 0.48 0.51 0.39 -61.22
Gold 0.50 0.51 0.19 0.00
GA 0.32 0.00 0.00 -13.04
GA 0.49 0.50 0.37 -61.22
GD 0.24 0.00 0.00 -17.72
GD 0.49 0.50 0.37 -62.18
IDK 0.48 0.51 0.30 -74.40
IDK 0.49 0.50 0.37 -63.15
KL 0.27 0.00 0.00 -13.92
KL 0.49 0.50 0.37 -62.18
NPO 0.49 0.51 0.36 -63.15
NPO 0.49 0.51 0.36 -64.13
Retain FT 0.49 0.51 0.36 -60.26
Retain FT 0.49 0.50 0.37 -61.22
RMU 0.27 0.00 0.00 -23.68
RMU 0.49 0.50 0.36 -61.22
LLMU 0.47 0.49 0.37 -73.34
LLMU 0.49 0.51 0.36 -60.26
DPO 0.46 0.49 0.39 -61.22
DPO 0.48 0.50 0.37 -65.12
SCRUB 0.49 0.51 0.36 -62.18
SCRUB 0.50 0.51 0.35 -61.22
Table 4: Results on experiments with and without LoRA regularization. The gray color shows that the method completely fails on the retain set.
6 Results and Discussion
Text domain
Table 1 presents the results for the text domain, showing that the RMU, KL, GD, and GA methods excel in unlearning the forget set (with the forget metric dropping to 0), but they suffer from catastrophic forgetting on the retain data (the retain metric also drops to 0). The remaining methods maintain performance on the retain set (retain metrics remain roughly the same), but their unlearning quality is poor – the forget metric is close to the one achieved by Retain FT. Identifying an optimal method that balances unlearning and retention is challenging. However, among the tested methods, IDK, DPO, and SCRUB provide the best (lowest) forget metrics without a drop in retain performance. These observations are consistent across both the LLaMa and Mistral models.
Visual MU results are presented in Table 2. It shows that most methods achieve high accuracy on the forget set with competitive U-LIRA and U-MIA values. Notably, $\text{SCRUB}_{bio}$ and Twins perform best across all considered metrics, making them optimal in this context. The Holdout Accuracy is relatively consistent across methods.
In multimodal unlearning, Table 3 shows that for the LLMU method, unlearning both modalities yields better results than text-only unlearning. The forget metric drops from 0.37 in the textual domain to 0.25 when unlearning both domains, while the retain and real metrics remain stable. For DPO, the results are less straightforward, but it is evident that unlearning the visual domain is crucial. When unlearning in the visual or both domains, the forget metric is 0.22, compared to 0.38 for text-only unlearning. The retain and real metrics stay consistent. However, SCRUB remains robust across all modalities, performing consistently in all three setups.
Then we run our experiments on unlearning both domains. The picture is very similar to the experiments on the textual domain only. Table 4 shows that GA, GD, KL and RMU effectively unlearn the forget set (forget metric goes to 0) but exhibit significant catastrophic forgetting of the retain set (retain metrics also goes to 0). In contrast, IDK, SCRUB, LLMU and DPO remain stable on the retain set (around 0.48), but their unlearning quality is worse (0.37 versus 0.39 on the original ununlearned model). The achieving the balance between unlearning and retention is challenging again.
Leaderboards We construct our leaderboards straightforwardly. First, we exclude methods that fail to retain knowledge from the retain set (highlighted in gray in the tables) and then rank the remaining methrods based on the Forget metric (or U-LIRA in the visual domain). The top-3 methods among each modality are shown in figure 2
LoRA Regularization Lastly, the experiments on $\ell_{1}$ LoRA regularization (Table 4) show improvements in unlearning quality for several methods, significantly reducing catastrophic forgetting in Gradient Ascent, Gradient Difference, KL minimization, RMU, and especially in LLMU. However, the proposed regularization is not beneficial for all unlearning techniques.
7 Conclusion
In this work, we introduce CLEAR, the first open-sourced benchmark designed to assess machine unlearning in multimodal (textual and visual) setup. Our evaluation of existing unlearning techniques across domains shows that multimodal unlearning is more challenging than previously anticipated, laying the ground for further research. Our studies on incorporating a LoRA regularization term demonstrate that this simple technique improves unlearning and can be easily integrated into other MU methods. We aim to encourage further research on enhancing privacy and security in large-scale AI models by offering an open-source benchmark. Future work could focus on improving MMU algorithms and expanding unlearning to new modalities, such as voice and video.
Limitations
Ethics
We utilized 84 hours of A100 GPU computation for our experiments, which resulted in an estimated 9 kg of CO2 emissions.
Appendix A Unlearning Methods
  1. 1.
  2. 2.
    $L(D_{F},\theta)=\frac{1}{|D_{F}|}\sum_{x\in D_{F}}{NLL}(x,\theta),$
  3. 3.
    $L_{GD}=-L(D_{F},\theta)+L(D_{R},\theta),$
  4. 4.
    The total objective function is formulated as follows:
    $L_{obj}=-L(D_{F},\theta)+L_{\text{KL}},$
  5. 5.
    $L_{idk}=L(D_{R},\theta)+L(D_{F}^{idk},\theta),$
  6. 6.
  7. 7.
    Negative Preference Optimization . Proposed in Zhang et al. (2024) this method can be treated as DPO without positive examples. In our setting, the final loss function $L_{NPO}$ for this method is derived as follows:
  8. 8.
    $d(x,w^{s})=\operatorname{KL}(p(f(x;w^{o}))||p(f(x;w^{s}))),$
    $L_{R}=\frac{\alpha}{|D_{R}|}\sum_{x_{r}\in D_{R}}d(x_{r},w^{s}),$
    $L_{F}=\frac{1}{|D_{F}|}\sum_{x_{f}\in D_{F}}d(x_{f},w^{s}),$
    $L_{\text{task}}=\frac{\gamma}{|D_{R}|}\sum_{x_{r}\in D_{R}}l(x_{r},y_{r}),$
    $L=L_{R}-L_{F}+L_{\text{task}},$
  9. 9.
    LLMU Yao et al. (2024b)
    This method was proposed in one of the first works on unlearning LLMs Yao et al. (2024b). In our experiments, we made slight modifications to the original method, and employed the following loss function:
    $L_{F}:=-L(D_{F},\theta),$
    $L_{LLMU}=L_{F}+L_{r}+L_{R},$
  10. 10.
    $L_{\operatorname{RMU}}=L_{\operatorname{F}}+L_{\operatorname{R}},$
  11. 11.
    $L_{\operatorname{F}}=d(f(x_{F}),f_{o}(x_{F}^{aug}))$
    $L_{\operatorname{R}}=d(f(x_{R}),f_{o}(x_{R})),$
    $L=L_{\operatorname{F}}+L_{R},$
  12. 12.
    $L=L_{\operatorname{F}}+L_{R},$
  13. 13.
    $L=L_{\operatorname{R}}+\lambda\cdot||\theta||_{1},$
    where $\lambda$ is a parameter of regularization.
  14. 14.
    Gradient Orhogonalization This method maximizes the loss of the original task on the forget set $D_{F}$ by ascending in the tangent direction of the gradient of the loss on the retain set $\nabla L_{R}$. The resulting weight update step is as follows:
    $\theta_{i+1}=\theta_{i}+\eta\left(\nabla L_{F}-\frac{(\nabla L_{F},\nabla L_{R})}{|\nabla L_{R}|}\nabla L_{R}\right)$
    where (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) is the scalar product, and $\eta$ is the learning rate. This method requires a very small learning rate and many unlearning epochs due to the instability and the complexity of convergence. In our experiments, we used 400 unlearning epochs followed by 100 epochs of finetuning. As shown in Figure 5, the effects of 400 unlearning epochs were effectively undone by just one epoch of subsequent finetuning on retain set $D_{R}$.
    Figure 5: Process of unlearning with tangent gradient maximization. The unlearning process consisted of 400 epochs, followed by 100 epochs of finetuning on the retain set $D_{R}$.
Appendix B The process of face generation
Appendix C A sample of dataset
@@
Appendix D Textual-only unlearning hyperparameters
For unlearning of the textual domain only, we use the textual part of CLEARbenchmark, containing question-answer pairs of about 200 authors, 20 for each of them (4000 pairs in total), and use the splits of size 90% and 10% of the entire data for retain and forget parts respectively. The "Gold" model for the further unlearning quality evaluation is trained on the retain data only, conducting 5 epochs of training with the batch size of 4, 1 gradient accumulation step, learning rate of $10^{-5}$, weight decay of 0.01, and also applying LoRA adapter with the rank 8, $\alpha=32$ and 0 dropout parameter. For the unlearning, we first finetune the model on the entire data split with the same hyperparameters: 5 epochs of training, batch size of 4, 1 gradient accumulation step, learning rate of $10^{-5}$, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, 0 dropout coefficient. Then, unlearning methods are conducted on the forget data with the following hyperparameters: 5 epochs of unlearning, batch size of 4, 1 gradient accumulation step, learning rate of $10^{-5}$, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, zero probability dropout. Such experimental settings and hyperparameters are the same for both Llama2-7B and Mistral architectures. To assess the unlearning quality, we compare the obtained unlearned model with the "gold" one and calculate ROUGE-L on retain and forget parts, Forget Quality and Model Utility metrics.
Appendix E CV pipeline
@@
For each of these 128 models, we run the forgetting algorithm on the forget subset of this particular model. From the resulting 128 models, we randomly select 64 target models (the remaining 64 will be used as shadow models for U-MIA and U-LIRA methods, see section G) on which the quality of the forgetting algorithms will be tested. Each of the 64 target models forgets a sample $\mathcal{D}f$ of 20 personalities. Additionally, for each target model, we form a holdout set $D_{H}$ by selecting 20 personalities that were not used in the training of this model.
@@
Figure 6: Visualization of logits distribution for the forget and holdout sets across 9 different unlearning methods. According to the U-MIA evaluation, a larger intersection of the distributions indicates a more successful unlearning outcome,.
Appendix F Multimodal unlearning hyperparameters
In a multimodal setting, we use both visual and textual parts of CLEARdataset, which consists of 4000 textual pairs of questions and answers about 200 authors, 20 for each of them, and 3770 images related to corresponding authors (number of images is less than the number of pairs because of GPT guard breaks and bugs in TOFU benchmark, as was described above). Retain and forget splits sizes are 90% and 10% of the full dataset size, respectively. The "Gold" model is trained on the retain data only with 3 epochs of training, batch size of 12, 1 gradient accumulation step, learning rate of $10^{-5}$, weight decay of 0.01, LoRA rank of 8, $\alpha=32$ and 0 dropout parameter. Unlearned models are also first finetuned on the full dataset with the same hyperparameters: 3 epochs of training, batch size of 12, 1 gradient accumulation step, learning rate of $10^{-5}$, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, 0 dropout parameter. After that, unlearning techniques are applied to the model on the forget data using the following hyperparameters: 5 epochs of unlearning, batch size of 1, 2 gradient accumulation steps, learning rate of $10^{-5}$, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, 0 dropout coefficient. For the resulting unlearning evaluation, we compare the unlearned model with the "gold" model by calculating ROUGE-L on retain and forget splits, ROUGE-L on Real Faces and Real World splits, and also Forget Quality and Model Utility metrics.
Appendix G U-MIA and U-LIRA

### v3

@@ Appendix C The process of face generation
Appendix D A sample of dataset
@@
Appendix E Forget Quality Metric
Maini et al. (2024) calculate a statistical test on the outputs of two models: an unlearned model and the gold model. The Truth Ratio metric is considered as output for its effectiveness in informativeness. To assess this metric, the Kolmogorov-Smirnov test is used to compare the distributions of Truth Ratios from both models. A high p-value suggests that the distributions are close, and so are unlearned and gold models; a low p-value indicates that distributions differ, and the unlearned model is far from gold.
Nevertheless, the application of statistical tests for model evaluation is uncommon and may be confusing; therefore, we conduct additional checks and compare it with common distribution distances, such as Jensen-Shannon and Wasserstein distances. We perform a simple experiment: take our dataset, randomly split it into 10 equal folds and train 10 models on the progressively larger subsets – starting with fold 1, then folds 1 to 2, and so on, up to folds 1 to 9, and finally all of the data. The latter model is considered as gold. We construct the Truth Ratios for each model and compare the resulting distributions with the gold model. The idea is, that the metric should be monotonic w.r.t. percent of data used in the train. The results are presented in Figure 5. We show that indeed, the p-value sometimes fails to represent the differences in the data. For example, the values for the 10 and 20 percents are equal. And the values for 60%, 80% are not monotonic. So, we consider to not using the p-value metric and move to JS distance.
Figure 5: Results of testing distances between distributions. JS stands for Jensen-Shannon distance, and WD – for Wasserstein distance. We show that unlike above metrics, the KS test p-value is not monotonic, which implies it may not be the best choice for Forget Quality metric.
Appendix F Textual-only Unlearning
For unlearning of the textual domain only, we use the TOFU benchmark, containing question-answer pairs of about 200 authors, 20 for each of them (4000 pairs in total), and use the splits of size 90% and 10% of the entire data for retain and forget parts respectively. The "Gold" model for the further unlearning quality evaluation is trained on the retain data only, conducting 5 epochs of training with the batch size of 4, 1 gradient accumulation step, learning rate of 1e-5 weight decay of 0.01, and also applying LoRA adapter with the rank 8, $\alpha=32$ and 0 dropout parameter. For the unlearning, we first finetune the model on the entire data split with the same hyperparameters: 5 epochs of training, batch size of 4, 1 gradient accumulation step, learning rate of 1e-5, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, 0 dropout coefficient. Then, unlearning methods are conducted on the forget data with the following hyperparameters: 5 epochs of unlearning, batch size of 4, 1 gradient accumulation step, learning rate of 1e-5, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, zero probability dropout. Such experimental settings and hyperparameters are the same for both Llama2-7B and Mistral architectures. To assess the unlearning quality, we compare the obtained unlearned model with the "gold" one and calculate ROUGE-L on retain and forget parts, Forget Quality and Model Utility metrics. The full results are available in table 4.
Appendix G Visual-only Unlearning
@@
For each of these 128 models, we run the forgetting algorithm on the forget subset of this particular model. From the resulting 128 models, we randomly select 64 target models (the remaining 64 will be used as shadow models for U-MIA and U-LIRA methods, see section I) on which the quality of the forgetting algorithms will be tested. Each of the 64 target models forgets a sample $\mathcal{D}f$ of 20 personalities. Additionally, for each target model, we form a holdout set $D_{H}$ by selecting 20 personalities that were not used in the training of this model.
The full results are available at table 5
@@
Figure 6: Visualization of logits distribution for the forget and holdout sets across 9 different unlearning methods. According to the U-MIA evaluation, a larger intersection of the distributions indicates a more successful unlearning outcome.
Appendix H Multimodal unlearning hyperparameters
In a multimodal setting, we use both visual and textual parts of CLEARdataset, which consists of 4000 textual pairs of questions and answers about 200 authors, 20 for each of them, and 3770 images related to corresponding authors (number of images is less than the number of pairs because of GPT guard breaks and bugs in TOFU benchmark, as was described above). Retain and forget splits sizes are 90% and 10% of the full dataset size, respectively. The "Gold" model is trained on the retain data only with 3 epochs of training, batch size of 12, 1 gradient accumulation step, learning rate of 1e-5, weight decay of 0.01, LoRA rank of 8, $\alpha=32$ and 0 dropout parameter. Unlearned models are also first finetuned on the full dataset with the same hyperparameters: 3 epochs of training, batch size of 12, 1 gradient accumulation step, learning rate of 1e-5, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, 0 dropout parameter. After that, unlearning techniques are applied to the model on the forget data using the following hyperparameters: 5 epochs of unlearning, batch size of 1, 2 gradient accumulation steps, learning rate of 1e-5, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, 0 dropout coefficient. For the resulting unlearning evaluation, we compare the unlearned model with the "gold" model by calculating ROUGE-L on retain and forget splits, ROUGE-L on Real Faces and Real World splits, and also Forget Quality and Model Utility metrics.
Appendix I U-MIA and U-LIRA